You invested $6000 between two accounts paying 4% and 8% annual interest. If the total interest earned for the year was $400, how much was invested at each rate?
Answer by mananth(16946)

Part I 4.00% per annum ------------- Amount invested =x
Part II 8.00% per annum ------------ Amount invested = y
6000
Interest----- 400.00

Part I 4.00% per annum ---x
Part II 8.00% per annum ---y
Total investment
x + 1 y= 6000 -------------1
Interest on both investments
4.00% x + 8.00% y= 400
Multiply by 100
4 x + 8 y= 40000.00 --------2
Multiply (1) by -4
we get
-4 x -4 y= -24000.00
Add this to (2)
0 x 4 y= 16000
divide by 4
y = 4000
Part I 4.00% $ 2000
Part II 8.00% $ 4000

CHECK
2000 --------- 4.00% ------- 80.00
4000 ------------- 8.00% ------- 320.00
Total -------------------- 400.00

m.ananth@hotmail.ca